OpenStudy (anonymous):

1/w+1/x divided by w^2-x^2 divided by wx

OpenStudy (callisto):

= (1/w+1/x ) / [(w^2-x^2) wx] = (w+x) / [(wx)^2 (w+x)(w-x)] = 1/ [(w-x)(wx)^2] =1/ (w^3 x^2-w^2 x^3)

OpenStudy (anonymous):

correct answer was 1/w-x

OpenStudy (anonymous):

thnks for your hlp

OpenStudy (callisto):

wait does your question means (1/w+1/x )divided by (w^2-x^2 divided by wx) or (1/w+1/x divided by w^2-x^2 )divided by wx ?

OpenStudy (callisto):

i thought it was the 2nd choice

OpenStudy (callisto):

if it is the first one, then = (1/w+1/x ) / [(w^2-x^2) /wx] = (w+x) (wx) / [(wx) (w+x)(w-x)] = 1/ (w-x)